OpenStudy (anonymous):

Solving system of 3 equations Two out of the 3 equations have 3 variables , one of the equations only has 2 variables. Help! I have no idea what to do! The problem is -6x+5y+z=12 -4x+3y=6 2x-4y+4z=-18

OpenStudy (anonymous):

The one with only 2 variables...add a 0z

OpenStudy (anonymous):

Take the first two equations. I'm going to eliminate the x values. 2(-6x+5y+z=12) -12x+10y+2z=24 -3(-4x+3y =6) 12x-9y =-18 y+2z=6 save this for later

OpenStudy (anonymous):

Now take the last two equations. I'll eliminate x againl -4x+3y =6 -4x+3y =6 2(2x-4y+4z=-18) 4x-8y+4z=-36 -5y+4z=-30 save this for later

OpenStudy (anonymous):

take the two saved equations and eliminate a variable...Im going to eliminate y y+2z=6 5(y+2z=6) 5y+10z=30 -5y+4z=-30 -5y+4z=-30 14z=0 z=0 Solve for x and y now.